two cars of equal mass(m1=m2=1000 kg) collide.the first car running at 80 km/h while the other car is running at 60 km/h.if hte cars got struck after the collison,with what velocity did the crashed cars move?


1
Expert's answer
2021-06-29T08:18:21-0400

Let's apply the law of conservation of momentum:


"m_1v_1-m_2v_2=(m_1+m_2)v_f,""v_f=\dfrac{m_1v_1-m_2v_2}{m_1+m_2},""v_f=\dfrac{1000\ kg\cdot22.2\ \dfrac{m}{s}-1000\ kg\cdot16.7\ \dfrac{m}{s}}{1000\ kg+1000\ kg}=2.75\ \dfrac{m}{s}."